What volume of 0.75 mol/L solution in a laboratory contains 0.20 mol of potassium chloride
1
Expert's answer
2020-09-30T04:17:14-0400

If there are 0.75 mol of KCl in 1 L the proportion may be written:

x=0.2×1/0.75=0.267(L)=267(ml).